OpenStudy (anonymous):

1.A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (anonymous):

A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air? : Let s = plane speed in still air then (s+40) = effective speed with the wind and (s-40) = effective speed against : The two trips are the same distance, write a distance equation Dist = speed * time 6(s-40) = 5(s+40) 6s - 240 = 5s + 200 6s - 5s = 200 + 240 s = 440 mph in still air : : Check this by finding the actual distances, they should be the same 6*400 = 2400 mi 5*480 = 2400 mi

OpenStudy (anonymous):

dous this work

OpenStudy (anonymous):

yea it does!

OpenStudy (anonymous):

k